# April 3, 2025

### v1

@@ 4. Existence of solutions for Problem $\mathcal{P}(\Omega)$
The formulation of this initial boundary value problem is inspired by the classical Signorini formulation (cf., e.g., [27]). We construct solutions of Problem $\mathcal{P}(\Omega)$ by studying its penalised version first. In what follows, we denoted by $\kappa>0$ a penalty parameter that is meant to approach zero. In what follows, we denote by $\{f\}^{-}$ the negative part of a function $f$, which is defined as $\{f\}^{-}:=-\min\{f,0\}$.
@@
that satisfies the following variational equations
$2\rho\,\int_{\Omega}\ddot{\bm{u}}_{\kappa}(t)\cdot\bm{v}\,\mathrm{d}x+\int_{\Omega}A^{ijk\ell}e_{k\|\ell}(\bm{u}_{\kappa}(t))e_{i\|j}(\bm{v})\,\mathrm{d}x\,\mathrm{d}t$
$\quad+\int_{\Omega}B^{ijk\ell}e_{k\|\ell}(\dot{\bm{u}}_{\kappa}(t))e_{i\|j}(\bm{v})\,\mathrm{d}x\,\mathrm{d}t$
$\quad-\dfrac{\chi_{(0,T_{0})}(t)}{\kappa}\int_{\Gamma}\{[\bm{I}+u_{i,\kappa}(t)\bm{e}^{i}]\cdot\bm{q}\}^{-}v_{i}\bm{e}^{i}\cdot\bm{q}\,\mathrm{d}\Gamma$
$=\int_{\Omega}f^{i}(t)v_{i}\,\mathrm{d}x\,\mathrm{d}t,$
for all $\bm{v}=(v_{i})\in\bm{V}(\Omega)$ in the sense of distributions in $(0,T)$, and that satisfies the following initial conditions
@@
where $\bm{u}_{0}=(u_{i,0})\in\bm{U}(\Omega)$, and $\bm{u}_{1}=(u_{i,1})\in\bm{V}(\Omega)$ are the same as in Problem $\mathcal{P}(\Omega)$. $\blacksquare$
@@
In the case treated in this paper, for a.a. $t\in(0,T)$, the displacement solving Problem $\mathcal{P}(\Omega)$ is, in general, of class $\bm{H}^{1}(\Omega)$, where $\Omega$ is a Lipschitz domain in $\mathbb{R}^{3}$.
Let us recall the following general result which can be found in, e.g., Lemma 3.1 in [36] or [16].
@@
Let $\mathscr{O}\subset\mathbb{R}^{m}$, with $m\geq 1$ an integer, be an open set. The operator {}$-\{\cdot\}^{-}:L^{2}(\mathscr{O})\to L^{2}(\mathscr{O})$ defined by
$f\in L^{2}(\mathscr{O})\mapsto-\{f\}^{-}:=\min\{f,0\}\in L^{2}(\mathscr{O}),$
@@
We re-cast the penalty term, which describes the energy that has to be paid by the system to violate the constraint, in terms of the non-linear operator $\mathcal{N}:\bm{V}(\Omega)\to\bm{V}^{\ast}(\Omega)$, that we define by:
@@
for all $\bm{w}=(w_{i})\in\bm{V}(\Omega)$ and all $\bm{v}=(v_{i})\in\bm{V}(\Omega)$, and we observe that the latter definition makes sense being the trace of an element in $\bm{V}(\Omega)$ of class $\bm{L}^{2}(\Gamma)$ when $\Omega$ is a Lipschitz domain as in our case. As a result, it is straightforward to observe that also the operator $\mathcal{N}$ is monotone, bounded and Lipschitz continuous.
We then define the non-linear operator $\tilde{\mathcal{N}}:L^{2}(0,T;\bm{V}(\Omega))\to L^{2}(0,T;\bm{V}^{\ast}(\Omega))$ pointwise by
(1) $(\tilde{\mathcal{N}}\bm{w})(t):={\mathcal{N}}(\bm{w}(t)),\quad\textup{ for a.a. }t\in(0,T),$
@@
We now recall the very important Gronwall inequality often used to study evolutionary problems: (see, e.g. the original paper [21] or, for instance, Theorem 1.1 in Chapter III of [23]).
@@
$\dfrac{\,\mathrm{d}y}{\,\mathrm{d}t}(t)\leq a(t)y(t)+b(t),\quad\textup{ for a.a. }t\in(0,T),$
where $a,b\in L^{1}(0,T)$ and $a(t),b(t)\geq 0$ for a.a. $t\in(0,T)$. Then, it results
@@
Assume that the applied body force $\bm{f}=(f_{i})\in L^{2}(0,T;\bm{L}^{2}(\Omega))$ is admissible in the sense of Definition 3.1. Then, Problem $\mathcal{P}_{\kappa}(\Omega)$ admits at least one solution.
@@
(i) Construction of a Galerkin approximation. Since $\bm{V}(\Omega)$ is an infinite dimensional separable Hilbert space which is also dense in $\bm{L}^{2}(\Omega)$, we infer that there exists an orthogonal basis $\{\bm{w}^{p}\}_{p=1}^{\infty}$ of the space $\bm{V}(\Omega)$, whose elements also constitute a Hilbert basis of the space $\bm{L}^{2}(\Omega)$.
@@
and satisfying the following variational equations
$2\rho\int_{\Omega}\ddot{\bm{u}}_{\kappa}^{m}(t)\cdot\bm{w}^{p}\,\mathrm{d}x+\int_{\Omega}A^{ijk\ell}e_{k\|\ell}(\bm{u}_{\kappa}^{m}(t))e_{i\|j}(\bm{w}^{p})\,\mathrm{d}x\,\mathrm{d}t$
$\quad+\int_{\Omega}B^{ijk\ell}e_{k\|\ell}(\dot{\bm{u}}_{\kappa}^{m}(t))e_{i\|j}(\bm{w}^{p})\,\mathrm{d}x\,\mathrm{d}t$
$\quad-\dfrac{\chi_{(0,T_{0})}(t)}{\kappa}\int_{\Gamma}\{[\bm{I}+u_{i,\kappa}^{m}(t)\bm{e}^{i}]\cdot\bm{q}\}^{-}w_{i}^{p}\bm{e}^{i}\cdot\bm{q}\,\mathrm{d}\Gamma$
$=\int_{\Omega}f^{i}(t)w_{i}^{p}\,\mathrm{d}x\,\mathrm{d}t,$
for all $1\leq p\leq m$ in the sense of distributions in $(0,T)$, and that satisfies the following initial conditions
@@
We observe that $\bm{u}_{0}^{m}\to\bm{u}_{0}$ in $\bm{V}(\Omega)$ as $m\to\infty$, and that $\bm{u}_{1}^{m}\to\bm{u}_{1}$ in $\bm{L}^{2}(\Omega)$ as $m\to\infty$. Since the coefficients $c_{p,\kappa}$ and their derivatives only depend on the time variable, we can take them outside of the integral sign, getting a $m\times m$ non-linear system of second order ordinary differential equations with respect to the variable $t$. Such a system can be rewritten in the form
@@
$\quad+\dfrac{\chi_{(0,T_{0})}(t)}{\kappa}\left(\int_{\Gamma}\left(\{[\bm{I}+(\bm{C}_{\kappa}(t)\cdot(w^{1}_{i}\dots w^{m}_{i}))\bm{e}^{i}]\cdot\bm{q}\}^{-}\right)(w^{p}_{i}\bm{e}^{i}\cdot\bm{q})\,\mathrm{d}x\right)_{p=1}^{m}$
$\quad+\left(\int_{\Omega}f^{i}(t)w_{i}^{p}\,\mathrm{d}x\right)_{p=1}^{m}$
@@
(3) $\rho\dfrac{\,\mathrm{d}}{\,\mathrm{d}t}\int_{\Omega}\dot{u}_{i,\kappa}^{m}(t)\dot{u}_{i}^{m}(t)\,\mathrm{d}x+\dfrac{1}{2}\dfrac{\,\mathrm{d}}{\,\mathrm{d}t}\int_{\Omega}A^{ijk\ell}e_{k\|\ell}(\bm{u}^{m}_{\kappa}(t))e_{i\|j}(\bm{u}^{m}_{\kappa}(t))\,\mathrm{d}x$
$\quad+\int_{\Omega}B^{ijk\ell}e_{k\|\ell}(\dot{\bm{u}}^{m}_{\kappa}(t))e_{i\|j}(\dot{\bm{u}}^{m}_{\kappa}(t))\,\mathrm{d}x$
$\quad+\dfrac{\chi_{(0,T_{0})}(t)}{2\kappa}\dfrac{\,\mathrm{d}}{\,\mathrm{d}t}\left(\int_{\Gamma}\left(\left\{[\bm{I}+u^{m}_{i,\kappa}(t)\bm{e}^{i}]\cdot\bm{q}\right\}^{-}\right)^{2}\,\mathrm{d}x\right)$
$=\int_{\Omega}f^{i}(t)\dot{u}^{m}_{i,\kappa}(t)\,\mathrm{d}x,$
@@
$\quad+\int_{0}^{t}\int_{\Omega}B^{ijk\ell}e_{k\|\ell}(\dot{\bm{u}}^{m}_{\kappa}(t))e_{i\|j}(\dot{\bm{u}}^{m}_{\kappa}(t))\,\mathrm{d}x\,\mathrm{d}t$
$\quad+\dfrac{\chi_{(0,T_{0})}(t)}{2\kappa}\int_{\Gamma}\left(\left\{[\bm{I}+u^{m}_{i,\kappa}(t)\bm{e}^{i}]\cdot\bm{q}\right\}^{-}\right)^{2}\,\mathrm{d}x$
$=\rho\int_{\Omega}u_{i,1}^{m}u_{i,1}^{m}\,\mathrm{d}x+\dfrac{1}{2}\int_{\Omega}A^{ijk\ell}e_{k\|\ell}(\bm{u}_{0}^{m})e_{i\|j}(\bm{u}_{0}^{m})\,\mathrm{d}x$
$\quad+\dfrac{\chi_{(0,T_{0})}(t)}{2\kappa}\int_{\Gamma}\left(\left\{[\bm{I}+u^{m}_{i,0}\bm{e}^{i}]\cdot\bm{q}\right\}^{-}\right)^{2}\,\mathrm{d}x$
@@
Let us observe that the third integral in the left hand side of (4) tends to zero, since $\bm{u}_{0}^{m}\to\bm{u}_{0}$ in $\bm{V}(\Omega)$ as $m\to\infty$, and $\bm{u}_{0}\in\bm{U}(\Omega)$. An application of the Cauchy-Schwarz inequality gives that:
(5) $\int_{0}^{t}\int_{\Omega}f^{i}(\tau)\dot{u}^{m}_{i,\kappa}(\tau)\,\mathrm{d}x\,\mathrm{d}\tau$
$\leq\left(\int_{0}^{T}\|\bm{f}(t)\|_{\bm{L}^{2}(\Omega)}^{2}\,\mathrm{d}t\right)^{1/2}\left(\int_{0}^{t}\|\dot{\bm{u}}^{m}_{\kappa}(\tau)\|_{\bm{L}^{2}(\Omega)}^{2}\,\mathrm{d}\tau\right)^{1/2}$

### v3

@@ 4. Existence of solutions for Problem $\mathcal{P}(\Omega)$
The formulation of this initial boundary value problem is inspired by the classical Signorini formulation (cf., e.g., [35]). We construct solutions of Problem $\mathcal{P}(\Omega)$ by studying its penalised version first. In what follows, we denoted by >0{{}}{}>0 a real penalty parameter that is meant to approach zero. In what follows, we denote by $\{f\}^{-}$ the negative part of a function $f$, which is defined as $\min\{f,0\}$.
@@
that satisfies the variational equations
$2\rho\,\int_{\Omega}\ddot{\bm{u}}_{\kappa}(t)\cdot\bm{v}\,\mathrm{d}x+\int_{\Omega}A^{ijk\ell}e_{k\|\ell}(\bm{u}_{\kappa}(t))e_{i\|j}(\bm{v})\,\mathrm{d}x\,\mathrm{d}t+\int_{\Omega}B^{ijk\ell}e_{k\|\ell}(\dot{\bm{u}}_{\kappa}(t))e_{i\|j}(\bm{v})\,\mathrm{d}x\,\mathrm{d}t$
$\quad-\dfrac{1}{\kappa}\int_{\Gamma}\{[\bm{I}+u_{i,\kappa}(t)\bm{e}^{i}]\cdot\bm{q}\}^{-}v_{i}\bm{e}^{i}\cdot\bm{q}\,\mathrm{d}\Gamma=\int_{\Omega}f^{i}(t)v_{i}\,\mathrm{d}x\,\mathrm{d}t,$
for all $\bm{v}=(v_{i})\in\bm{V}(\Omega)$ in the sense of distributions in $(0,T)$, and that satisfies the initial conditions
@@
where $\bm{u}_{0}=(u_{i,0})\in\bm{U}(\Omega)$, and $\bm{u}_{1}=(u_{i,1})\in\bm{L}^{2}(\Omega)$ are the same as in Problem $\mathcal{P}(\Omega)$. $\blacksquare$
@@
In the case treated in this paper, for a.a. $t\in(0,T)$, the displacement solving Problem $\mathcal{P}(\Omega)$ is, in general, of class $\bm{H}^{1}(\Omega)$, where $\Omega$ is a Lipschitz domain in $\mathbb{R}^{3}$. Let us recall the following general result which can be found in, e.g., Lemma 3.1 in [46] or [1, 22].
@@
Let $\mathscr{O}\subset\mathbb{R}^{m}$, with $m\geq 1$ an integer, be a non-empty open set. The operator $\{\cdot\}^{-}:L^{2}(\mathscr{O})\to L^{2}(\mathscr{O})$ defined by
$L^{2}(\mathscr{O})\ni f\mapsto-\{f\}^{-}:=\min\{f,0\}\in L^{2}(\mathscr{O}),$
@@
We re-cast the penalty term, which describes the energy that has to be paid by the system to violate the constraint, in terms of the non-linear operator $\bm{V}(\Omega)\to\bm{V}^{\ast}(\Omega)$, that we define by $\mathcal{N}:=\left(\textup{tr}^{\star}\circ(-\{\cdot\}^{-})\circ\textup{tr}\right)$, where is the Banach adjoint of tr. The operator $\mathcal{N}$ is such that
@@
for all $\bm{w}=(w_{i})\in\bm{V}(\Omega)$ and all $\bm{v}=(v_{i})\in\bm{V}(\Omega)$. Observe that the latter definition makes sense being the trace of an element in $\bm{V}(\Omega)$ of class $\bm{L}^{2}(\Gamma)$ when $\Omega$ is a Lipschitz domain as in our case. As a result, it is straightforward to observe that also the operator $\mathcal{N}$ is monotone, bounded and Lipschitz continuous.
We then define the non-linear operator $L^{2}(0,T;\bm{V}(\Omega))\to L^{2}(0,T;\bm{V}^{\ast}(\Omega))$ pointwise by
[…]
@@
We now recall the very important Gronwall’s inequality often used to study evolutionary problems (see, e.g. the original paper [28] or, for instance, Appendix B.2(j) in [21] or Theorem 1.1 in Chapter III of [30]).
@@
$\dot{y}(t)\leq a(t)y(t)+b(t),\quad\textup{ for a.a. }t\in(0,T),$
where $a,b\in L^{1}(0,T)$ and $a(t),b(t)\geq 0$ for a.a. $t\in(0,T)$. Then, it results:
@@
Let us also recall a classical result in the analysis of evolutionary equations: the Aubin-Lions-Simon theorem (cf., e.g., Theorem 8.62 in [37]).
Theorem 4.2 (Aubin-Lions-Simon).
Let $I\subset\mathbb{R}$ be an open and bounded interval. Let $(Y_{0},\|\cdot\|_{Y_{0}})$, $(Y_{1},\|\cdot\|_{Y_{1}})$ and $(Y_{2},\|\cdot\|_{Y_{2}})$ be Banach spaces such that:
$Y_{0}\hookrightarrow\hookrightarrow Y_{1}\hookrightarrow Y_{2}.$
Let $1\leq p<\infty$ and let $1\leq q\leq\infty$, and let $\mathcal{V}$ be the Banach space defined by $\mathcal{V}:=\{u\in L^{p}(I;Y_{0});\dot{u}\in L^{q}(I;Y_{2})\}$, and equipped with the norm:
[…]
Then, the continuous embedding $\mathcal{V}\hookrightarrow L^{p}(I;Y_{1})$ is compact. If $p=\infty$ and $q>1$ then the continuous embedding $\mathcal{V}\hookrightarrow\mathcal{C}^{0}(\overline{I};Y_{1})$ is compact. ∎
The next step consists in showing, by Galerkin method, that Problem $\mathcal{P}_{\kappa}(\Omega)$ admits at least one solution. Prior to establishing this result, we need a lemma on the compactness of the time-dependent version of the trace operator. We establish this lemma in a general setting.
Lemma 4.2.
Let $\bm{\mathcal{V}}:=\{\bm{v}\in L^{2}(0,T;\bm{V}(\Omega));\dot{\bm{v}}\in L^{2}(0,T;\bm{V}^{\ast}(\Omega))\}$. Let $\textup{tr}:\bm{V}(\Omega)\to\bm{L}^{2}(\Gamma)$ denote the classical trace operator. Then the operator $\tilde{\textup{tr}}:\bm{\mathcal{V}}\to L^{2}(0,T;\bm{L}^{2}(\Gamma))$, that constitutes the time-dependent version of tr, and that is defined pointwise by
[…]
at each $\bm{v}\in\bm{\mathcal{V}}$ is linear, continuous and compact.
Proof.
The linearity and continuity properties are straightforward, and the proof closely follows the strategy for establishing the linearity and continuity of the time-dependent version of the linearised change of metric tensor (cf., e.g., page 4 in [44]).
To establish the compactness, consider a sequence $\{\bm{v}_{n}\}_{n=1}^{\infty}$ that is bounded in $\bm{\mathcal{V}}$. An application of the Aubin-Lions-Simon theorem (Theorem 4.2) shows that $\bm{v}_{n}\to\bm{v}$ in $L^{2}(0,T;\bm{L}^{2}(\Omega))$ as $n\to\infty$ up to passing to a subsequence. Thanks to Theorem 18.1(iii) and Corollary 18.4 in [37], there exits two constants $c=c(\Omega)>0$ and $\epsilon=\epsilon(\Omega)>0$ which are independent of $n$ and $t$ such that:
(5) $\int_{\Gamma}|\textup{tr}(\bm{v}_{n}(t))-\textup{tr}(\bm{v}(t))|^{2}\,\mathrm{d}\Gamma\leq\dfrac{c}{\varepsilon}\int_{\Omega}|\bm{v}_{n}(t)-\bm{v}(t)|^{2}\,\mathrm{d}x+c\varepsilon\sum_{i=1}^{3}\int_{\Omega}|\nabla v_{i,n}(t)-\nabla v_{i}(t)|^{2}\,\mathrm{d}x,$
for all $0<\varepsilon<\epsilon$ and for a.a. $t\in(0,T)$. Integrating (5) in $(0,T)$ gives:
(6) $\int_{0}^{T}\int_{\Gamma}|(\tilde{\textup{tr}}\,\bm{v}_{n})(t)-(\tilde{\textup{tr}}\,\bm{v})(t)|^{2}\,\mathrm{d}\Gamma\,\mathrm{d}t=\int_{0}^{T}\int_{\Gamma}|\textup{tr}(\bm{v}_{n}(t))-\textup{tr}(\bm{v}(t))|^{2}\,\mathrm{d}\Gamma\,\mathrm{d}t$
$\leq\dfrac{c}{\varepsilon}\int_{0}^{T}\int_{\Omega}|\bm{v}_{n}(t)-\bm{v}(t)|^{2}\,\mathrm{d}x\,\mathrm{d}t+c\varepsilon\sum_{i=1}^{3}\int_{0}^{T}\int_{\Omega}|\nabla v_{i,n}(t)-\nabla v_{i}(t)|^{2}\,\mathrm{d}x\,\mathrm{d}t.$
Consequently, the first term on the right-hand side of the inequality in (6) tends to zero as $n\to\infty$.
Additionally, thanks to the assumed boundedness for the sequence $\{\bm{v}_{n}\}_{n=1}^{\infty}$ in $L^{2}(0,T;\bm{V}(\Omega))$, the second term on the right-hand side of the inequality in (6) is bounded independently of $n$. Therefore, there exists a constant $\tilde{C}>0$ independent of $t$ and $n$ such that:
(7) $\limsup_{n\to\infty}\int_{0}^{T}\int_{\Gamma}|(\tilde{\textup{tr}}\,\bm{v}_{n})(t)-(\tilde{\textup{tr}}\bm{v})(t)|^{2}\,\mathrm{d}\Gamma\,\mathrm{d}t\leq\tilde{C}\varepsilon.$
Thanks to the arbitrariness of $0<\varepsilon\leq\epsilon$, we obtain that:
$\lim_{n\to\infty}\int_{0}^{T}\int_{\Gamma}|(\tilde{\textup{tr}}\,\bm{v}_{n})(t)-(\tilde{\textup{tr}}\bm{v})(t)|^{2}\,\mathrm{d}\Gamma\,\mathrm{d}t=0,$
thus showing that $\tilde{\textup{tr}}\,\bm{v}_{n}\to\tilde{\textup{tr}}\,\bm{v}$ in $L^{2}(0,T;\bm{L}^{2}(\Gamma))$ as $n\to\infty$, and establishing the sought compactness. Note that the subsequence for which the pre-compactness of $\{\tilde{\textup{tr}}\,\bm{v}_{n}\}_{n=1}^{\infty}$ is realised is exactly the one for which the convergence $\bm{v}_{n}\to\bm{v}$ in $L^{2}(0,T;\bm{L}^{2}(\Omega))$ as $n\to\infty$ asserted by the Aubin-Lions-Simon theorem holds. ∎
Next, we establish the continuity of the operators $\mathcal{A}$ and $\mathcal{B}$.
Lemma 4.3.
The operator $\mathcal{A}:\bm{V}(\Omega)\to\bm{V}^{\ast}(\Omega)$ defined in (1) is linear and continuous. Moreover, also the corresponding time-dependent version $L^{\infty}(0,T;\bm{V}(\Omega))\to L^{\infty}(0,T;\bm{V}^{\ast}(\Omega))$ defined pointwise by
[…]
for a.a. $t\in(0,T)$ is linear and continuous.
Proof.
Lemma 4.4.
The operator $\mathcal{B}:\bm{V}(\Omega)\to\bm{V}^{\ast}(\Omega)$ defined in (2) is linear and continuous. Moreover, also the corresponding time-dependent version $L^{2}(0,T;\bm{V}(\Omega))\to L^{2}(0,T;\bm{V}^{\ast}(\Omega))$ defined pointwise by
[…]
for a.a. $t\in(0,T)$ is linear and continuous. ∎
We are now ready to establish the existence of solutions for Problem $\mathcal{P}_{\kappa}(\Omega)$ via Galerkin method. We note in passing that the presence of the non-linear term associated with the extent to which the constraint is broken renders the problem challenging. The proof of the next result hinges on the compactness of the time-dependent version of the trace operator established in Lemma 4.2.
@@
Assume that $\bm{f}=(f^{i})\in L^{2}(0,T;\bm{L}^{2}(\Omega))$. Then Problem $\mathcal{P}_{\kappa}(\Omega)$ admits at least one solution.
@@
(i) Construction of a Galerkin approximation. Since $\bm{V}(\Omega)$ is an infinite dimensional separable Hilbert space which is dense in $\bm{L}^{2}(\Omega)$ and compactly embedded in $\bm{L}^{2}(\Omega)$ in light of the Rellich-Kondrašov theorem (cf., e.g., Theorem 8.4-3 in [15]), we infer that there exists an orthogonal basis $\{\bm{w}^{p}\}_{p=1}^{\infty}$ of the space $\bm{V}(\Omega)$, whose elements also constitute a Hilbert basis of the space $\bm{L}^{2}(\Omega)$.
@@
and satisfying the variational equations
$2\rho\int_{\Omega}\ddot{\bm{u}}_{\kappa}^{m}(t)\cdot\bm{w}^{p}\,\mathrm{d}x+\int_{\Omega}A^{ijk\ell}e_{k\|\ell}(\bm{u}_{\kappa}^{m}(t))e_{i\|j}(\bm{w}^{p})\,\mathrm{d}x+\int_{\Omega}B^{ijk\ell}e_{k\|\ell}(\dot{\bm{u}}_{\kappa}^{m}(t))e_{i\|j}(\bm{w}^{p})\,\mathrm{d}x$
$\quad-\dfrac{1}{\kappa}\int_{\Gamma}\{[\bm{I}+u_{i,\kappa}^{m}(t)\bm{e}^{i}]\cdot\bm{q}\}^{-}w_{i}^{p}\bm{e}^{i}\cdot\bm{q}\,\mathrm{d}\Gamma=\int_{\Omega}f^{i}(t)w_{i}^{p}\,\mathrm{d}x,$
for all $1\leq p\leq m$ in the sense of distributions in $(0,T)$, and that satisfies the initial conditions
@@
We observe (cf., e.g., Theorem 4.13-1 in [15]) that $\bm{u}_{0}^{m}\to\bm{u}_{0}$ in $\bm{V}(\Omega)$ as $m\to\infty$, and that $\bm{u}_{1}^{m}\to\bm{u}_{1}$ in $\bm{L}^{2}(\Omega)$ as $m\to\infty$. Since the coefficients $c_{p,\kappa}$ and their derivatives only depend on the time variable, we can take them outside of the integral sign, getting a $m\times m$ non-linear system of second order ordinary differential equations with respect to the variable $t$. Such a system can be rewritten in the form
@@
$\quad+\dfrac{1}{\kappa}\left(\int_{\Gamma}\left(\{[\bm{I}+(\bm{C}_{\kappa}(t)\cdot(w^{1}_{i}\dots w^{m}_{i}))\bm{e}^{i}]\cdot\bm{q}\}^{-}\right)(w^{p}_{i}\bm{e}^{i}\cdot\bm{q})\,\mathrm{d}x\right)_{p=1}^{m}$
$\quad+\left(\int_{\Omega}f^{i}(t)w_{i}^{p}\,\mathrm{d}x\right)_{p=1}^{m},$
@@
(9) $\rho\dfrac{\,\mathrm{d}}{\,\mathrm{d}t}\int_{\Omega}\dot{u}_{i,\kappa}^{m}(t)\dot{u}_{i,\kappa}^{m}(t)\,\mathrm{d}x+\dfrac{1}{2}\dfrac{\,\mathrm{d}}{\,\mathrm{d}t}\int_{\Omega}A^{ijk\ell}e_{k\|\ell}(\bm{u}^{m}_{\kappa}(t))e_{i\|j}(\bm{u}^{m}_{\kappa}(t))\,\mathrm{d}x+\int_{\Omega}B^{ijk\ell}e_{k\|\ell}(\dot{\bm{u}}^{m}_{\kappa}(t))e_{i\|j}(\dot{\bm{u}}^{m}_{\kappa}(t))\,\mathrm{d}x$
$\quad+\dfrac{1}{2\kappa}\dfrac{\,\mathrm{d}}{\,\mathrm{d}t}\left(\int_{\Gamma}\left(\left\{[\bm{I}+u^{m}_{i,\kappa}(t)\bm{e}^{i}]\cdot\bm{q}\right\}^{-}\right)^{2}\,\mathrm{d}\Gamma\right)=\int_{\Omega}f^{i}(t)\dot{u}^{m}_{i,\kappa}(t)\,\mathrm{d}x,$
@@
$\quad+\int_{0}^{t}\int_{\Omega}B^{ijk\ell}e_{k\|\ell}(\dot{\bm{u}}^{m}_{\kappa}(\tau))e_{i\|j}(\dot{\bm{u}}^{m}_{\kappa}(\tau))\,\mathrm{d}x\,\mathrm{d}\tau+\dfrac{1}{2\kappa}\int_{\Gamma}\left(\left\{[\bm{I}+u^{m}_{i,\kappa}(t)\bm{e}^{i}]\cdot\bm{q}\right\}^{-}\right)^{2}\,\mathrm{d}x$
$=\rho\int_{\Omega}u_{i,1}^{m}u_{i,1}^{m}\,\mathrm{d}x+\dfrac{1}{2}\int_{\Omega}A^{ijk\ell}e_{k\|\ell}(\bm{u}_{0}^{m})e_{i\|j}(\bm{u}_{0}^{m})\,\mathrm{d}x+\dfrac{1}{2\kappa}\int_{\Gamma}\left(\left\{[\bm{I}+u^{m}_{i,0}\bm{e}^{i}]\cdot\bm{q}\right\}^{-}\right)^{2}\,\mathrm{d}x$
@@
Let us observe that the third integral in the right-hand side of (10) tends to zero, since $\bm{u}_{0}^{m}\to\bm{u}_{0}$ in $\bm{V}(\Omega)$ as $m\to\infty$, and $\bm{u}_{0}\in\bm{U}(\Omega)$ by assumption. An application of the Cauchy-Schwarz inequality gives:
(11) $\int_{0}^{t}\int_{\Omega}f^{i}(\tau)\dot{u}^{m}_{i,\kappa}(\tau)\,\mathrm{d}x\,\mathrm{d}\tau\leq\left(\int_{0}^{T}\|\bm{f}(t)\|_{\bm{L}^{2}(\Omega)}^{2}\,\mathrm{d}t\right)^{1/2}\left(\int_{0}^{t}\|\dot{\bm{u}}^{m}_{\kappa}(\tau)\|_{\bm{L}^{2}(\Omega)}^{2}\,\mathrm{d}\tau\right)^{1/2}$